OpenStudy (anonymous):

How come, when you continuosly press the sqrt sign on a calculator for a number, you get 1?

OpenStudy (anonymous):

Is it just the amount of data a calculator can hold for such a small number?

OpenStudy (anonymous):

depends on what number you start with

OpenStudy (anonymous):

if you start with a number bigger than one you should get 1

OpenStudy (anonymous):

Sorry, I meant 1, lol

OpenStudy (anonymous):

cause anything raised to the power 1 is just 1...

OpenStudy (anonymous):

\[1^2 = 1\]

OpenStudy (anonymous):

\[\sqrt{1} = 1\]

OpenStudy (anonymous):

@satellite73 not necessary to be greater than 1. any positive number will eventually reach 1

OpenStudy (anonymous):

true

OpenStudy (anonymous):

because your calculator round off the answer.

OpenStudy (anonymous):

not necessary to be greater than 1. any positive number will eventually reach 1

OpenStudy (anonymous):

\(1 \times 1 = 1\) or \(1^2 = 1\) or \(\sqrt{1} = 1\) No matter how many times you take a square root of 1 the answer will be 1. Anything to a power of 1 = itself. Anything to a power of 0 = 1.

OpenStudy (anonymous):

Makes sense